OpenStudy (anonymous):

i am having a problem for this question got 3 different answers could someone show me the correct formula? 12000 invested 6%r compounded semiannually thanks in advance

OpenStudy (anonymous):

FV = PV (1 + .03)^2 FV = 12,000 (1.03^2) FV = 12,000 (1.0609) FV = 12,730.8 = 12,731 Did you get this ?

OpenStudy (anonymous):

check this out... http://www.accountingcoach.com/online-accounting-course/84Xpg02.html

OpenStudy (anonymous):

thank you guys sorry this typing box froze on me

OpenStudy (anonymous):

not a problem....my typing box freezes sometimes....its annoying huh?